OpenStudy (anonymous):

Find and simplify dy/dx for yx-2y=x+2. Text book answer is -4/(x-2)^2 but that is not the answer I am coming up with.

OpenStudy (campbell_st):

start by making y the subject y(x - 2) = x + 2 now divide both sides by (x -2) \[y = \frac{x + 2}{x -2} \] you now need the quotient rule to find the derivative

OpenStudy (anonymous):

Thank you!